the combustion of methane produces carbon dioxide and water. assume that 2.5 mol of ch4 burned in the presence of excess air. what is the percent yield if the reaction produces 77.0g of CO2?
1
Expert's answer
2020-04-10T13:35:49-0400

CH4+2O2=CO2+2H2O


n(CO2):n(CH4)=1:1=2.5mol

m(CO2)=n(CO2)*M(CO2)=2.5*44=110g


Ƞ=(mpr/mt)*100%=(77/110)/100%=70%


Answer: Ƞ=70%